OpenStudy (anonymous):

What is half of 5 1/4?

jimthompson5910 (jim_thompson5910):

5 1/4 = 5+1/4 = 20/4+1/4 = (20+1)/4 = 21/4 Now divide 21/4 by 2 to get \[\frac{\frac{21}{4}}{2}=\frac{21}{4}\frac{1}{2}=\frac{21*1}{4*2}=\frac{21}{8}\] So \[\frac{\frac{21}{4}}{2}=\frac{21}{8}\] which means that half of 5 1/4 is \[\frac{21}{8}\] which is the mixed number 2 5/8

OpenStudy (anonymous):

thks